ganeshie8 (ganeshie8):

Suppose \(g(n) = \sum\limits_{i=1}^n f(i)\) then \(f(n) = ?\)

ganeshie8 (ganeshie8):

@ikram002p @imqwerty

OpenStudy (ikram002p):

g(n) is multiplicative or f(n) ?

ganeshie8 (ganeshie8):

g(n) is just a partial sum g or f need not be multiplicative

ganeshie8 (ganeshie8):

Suppose \(g(n) = \sum\limits_{i=1}^n f(i)\) then \(f(n) = ?\)

ganeshie8 (ganeshie8):

basically we want to express \(f\) as a function of \(g\)

ganeshie8 (ganeshie8):

in other words, we want the \function, \(f\), back

OpenStudy (ikram002p):

ok g(n)=f(1)+f(2)+...+f(n) g(n)-f(n)=f(1)+...+f(n-1) f(n)=g(n-1)-g(n)

ganeshie8 (ganeshie8):

Perfect!

ganeshie8 (ganeshie8):

wait, there seems to be a sign error... check again

imqwerty (imqwerty):

:o this is how we do it :) thanks!! :D

OpenStudy (ikram002p):

wait no i have made some mistake f(n)=g(n)-g(n-1) :P

ganeshie8 (ganeshie8):

Yes :) its a trivial problem... but I think it is a very good problem to enter the beautiful world of mobius inversion and stuff!

OpenStudy (ikram002p):

its recursive formula ok keep going

ganeshie8 (ganeshie8):

In light of above result, try finding below : Suppose \(g(n) = \sum\limits_{d\mid n} f(d)\) then \(f(n) = ?\)

OpenStudy (ikram002p):

then i need a shower lol

ganeshie8 (ganeshie8):

its not going to be trivial

ganeshie8 (ganeshie8):

you're euler if you're able to figure out \(f(n)\) on your own!

ganeshie8 (ganeshie8):

shower may not help that much..

ganeshie8 (ganeshie8):

here is the complete theorem :

OpenStudy (ikram002p):

well i start to think of it as this. if \( n=p_1p_2...p_k\) \(g(n)=\sum_{d|n} f(n)= f(p_1)f(p_2)...f(p_k)\)

ganeshie8 (ganeshie8):

Theorem : If \(g(n) = \sum\limits_{d\mid n} f(d)\) then \(f(n)\) is given by \[f(n) = \sum\limits_{d\mid n}\mu(d)g(n/d) \]

OpenStudy (ikram002p):

oh i was so close to that :P

imqwerty (imqwerty):

i ws thinking like how we solved the previous problem if we do that we get this-\[g(n)-\left[ f(d_1)+f(d_2)+f(d_3).. \right]=f(n)\] so we need some g(n) which must equal \(f(d_1)+f(d_2)+f(d_3)..\) so as to express f(n) in terms of g(n)

ganeshie8 (ganeshie8):

here is an interesting result based on that theorem : we know that \(\tau(n) = \sum\limits_{d\mid n}1\) applying previous theorem gives \(1 = \sum\limits_{d\mid n} \mu(d)\tau(n/d)\)

OpenStudy (ikram002p):

impressive! ok brb

ganeshie8 (ganeshie8):

that is a nice start !

ganeshie8 (ganeshie8):

it took me several weeks to wrap my head around this theorem... I would be amazed if somebody gets this on the first day itself !

ganeshie8 (ganeshie8):

ikram, f and g need not be multiplicative...

imqwerty (imqwerty):

lol m getting dizzy now xD i'll be back after i finish my home wrk :)

OpenStudy (ikram002p):

ermm i know, that property i used earlier which seems i'm the only one who use it xD

ganeshie8 (ganeshie8):

we had a lot of number theory for one day haha maybe lets get back on this tomorrow

OpenStudy (ikram002p):

sure!

OpenStudy (ikram002p):

ermmm lol i hope it won't take me that much.